Solve and simplify with an integer of fraction.
2/z = 7/z - 1/2
Thank you so much!!!

Answer by checkley77(12844)

2/z = 7/z - 1/2
2/z-7/z=-1/2
(2-7)/z=-1/2
5/z=-1/2 cross multiply.
-z=2*5
-z=10
z=-10 ans.
Proof:
2/10=7/10+1/2
2/10=7/10-5/10
2/10=2/10